OpenStudy (anonymous):

Find the lateral area of a right circular cone whose radius is 6 feet and height is 8 feet.

OpenStudy (anonymous):

Formula \[A _{L}=\pi r \sqrt{h ^{2} + r ^{2}} \]

OpenStudy (anonymous):

\[\pi*6*\sqrt{8^{2}+6^{2}}\]= Lateral Area

OpenStudy (anonymous):

is it 60π ft2